OpenStudy (anonymous):

Write the definite integral for the summation: the limit as n goes to infinity of the summation from k equals 1 to n of the product of the square of the quantity 1 plus k over n squared and 1 over n.

OpenStudy (anonymous):

https://gyazo.com/478f27e7e70fd5a27beb5d16043fe87e

OpenStudy (anonymous):

As n approaches to infinity the values get closer to 0

Miracrown (miracrown):

Indeed, As n approaches to infinity the values get closer to 0

OpenStudy (anonymous):

Sorry went to use the restroom.

OpenStudy (anonymous):

And now that I take a good look what I was thinking would make no sense here.

jimthompson5910 (jim_thompson5910):

Hint: \[\Large \lim_{n \to \infty}\sum_{k = 1}^{n}\left(1+\frac{k}{n}\right)^2\left(\frac{1}{n}\right)\] \[\Large \lim_{n \to \infty}\sum_{k = 1}^{n}\left(1+\frac{1}{n}*k\right)^2\left(\frac{1}{n}\right)\] \[\Large \lim_{n \to \infty}\sum_{k = 1}^{n}\left({\color{red}{1}}+{\color{blue}{\frac{1}{n}}}*k\right)^2\left({\color{blue}{\frac{1}{n}}}\right)\] \[\Large \lim_{n \to \infty}\sum_{k = 1}^{n}\left({\color{red}{a}}+{\color{blue}{\Delta x}}*k\right)^2\left({\color{blue}{\Delta x}}\right)\]

jimthompson5910 (jim_thompson5910):

Keep in mind that \[\Large x_k = a+\Delta x*k\] \[\Large \Delta x = \frac{b-a}{n}\]

OpenStudy (anonymous):

So b=2

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

But what is Xk for?

jimthompson5910 (jim_thompson5910):

xk represents a general term in the sequence x1,x2,x3, ..., xn

jimthompson5910 (jim_thompson5910):

the idea is that you can find the approx area under the curve by breaking up the interval into n equal pieces and finding the area of each rectangular piece

jimthompson5910 (jim_thompson5910):

this might help visualize an example https://www.math.hmc.edu/calculus/tutorials/riemann_sums/gif/figure6.gif

OpenStudy (anonymous):

But Xk would be equal to what's in the first parenthesis so how will that help me when writing the integral?

OpenStudy (anonymous):

Integral is equal to area under curve

jimthompson5910 (jim_thompson5910):

Hint: f(x) ----> f(xk)

OpenStudy (anonymous):

What?

jimthompson5910 (jim_thompson5910):

Do you see how I got\[\Large \lim_{n \to \infty}\sum_{k = 1}^{n}\left({\color{red}{1}}+{\color{blue}{\frac{1}{n}}}*k\right)^2\left({\color{blue}{\frac{1}{n}}}\right)\]

OpenStudy (anonymous):

Yes I understand the concept just not how it will be applied

jimthompson5910 (jim_thompson5910):

that will turn into \[\Large \lim_{n \to \infty}\sum_{k = 1}^{n}\left(x_k\right)^2\left({\color{black}{\frac{1}{n}}}\right)\]

jimthompson5910 (jim_thompson5910):

the \(\LARGE (x_k)^2\) portion represents the f(x) function the remaining bit is the delta x

OpenStudy (anonymous):

Ohhhhhhhh I feel stupid now :/

OpenStudy (anonymous):

Thanks

Miracrown (miracrown):

Don't feel stupid. You're still learning :)

jimthompson5910 (jim_thompson5910):

we have n rectangles each rectangle is delta x = 1/n units wide the height of each rectangle is equal to f(x_k) = (1+k/n)^2 units

OpenStudy (anonymous):

Got it Thanks